Circular No. 5990

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

4U 0115+634 D. M. Scott, M. H. Finger, R. B. Wilson, and T. A. Prince report

for the Compton Gamma Ray Observatory BATSE pulsar team: “Pulsed hard x-ray flux in the energy range 20–70 keV is currently being detected from the x-ray binary 4U 0115+634. This is the second outburst of this source observed by BATSE, the first occurring in April 1991. In the current outburst, pulsed flux was first detected on 1994 May 8. An intrinsic pulse period of 3.6145107 ± 0.0000010 s and a period derivative of (5.9 ± 1.9) × 10^{−11} were determined for epoch May 13.0 UT. The phase-averaged pulsed flux, averaged over 20–40 keV, was ∼ 45 mCrab (total) on May 15. The spectrum is well fit by an optically thin thermal bremsstrahlung model A/E exp(−kT), with kT = 16 ± 4 keV, and a flux at 35 keV of (2.0 ± 0.2) × 10^{−4} photon cm^{−2} s^{−1} keV^{−1}.”

SUPERNOVA 1994I IN NGC 5194 T. Tongue, M. Graham, M. Stanley, L. Truesdell, and D. Westpfahl,

New Mexico Institute of Mining and Technology, report a pre-supernova radio continuum observation of NGC 5194 using the Very Large Array on Feb. 12 (49 days before the first optical detection): “Using the continuum channels of a 21-cm spectral line observation, with baselines out to 50 000 wavelengths, we report a nondetection with a 3σ upper limit of 2.3 mJy at the position of the supernova. The upper limit is based on both random and systematic (calibration) errors; the rms noise of the continuum image was 0.75 mJy/beam. A nondetection at this level suggests that the progenitor star did not produce bright radio continuum emission from pre-explosion ejecta on timescales of a few months prior to the explosion and was not associated with a radio-bright H II region.”

CCD photometry by H. Mikuž, Ljubljana, Slovenia (cf. IAUC 5972): May 3.99 UT, V = 14.43 ± 0.03, B − V = +0.69 ± 0.05.

OJ 287 H. Meusinger, Thüringer Landessternwarte, Tautenburg, reports the

following photographic B magnitudes of OJ 287 (cf. IAUC 5909) from Tautenburg Schmidt plates (emulsion ORWO ZU, filter GG13): Feb. 14.9 UT, 16.1; Mar. 10.9, 16.0; Apr. 5.8, 15.2; 14.9, 15.75; May 2.9, 16.0; 7.9, 15.3; 9.9, 15.6; 13.9, 15.7.

1994 May 18 Daniel W. E. Green